[figure: 9.7 cm]

Competition between superconductivity and the pseudogap phase in

the t-J model

R. Zeyher and A. Greco^{a} Max-Planck-Institut für Festkörperforschung, 70569 Stuttgart, Germany aPermanent address: Departamento de Física, Facultad de Ciencias Exactas e Ingeniería, 2000-Rosario, Argentina

Extending the spin components from 2 to N and using 1/N as an expansion parameter the phase diagram of the t-J model on a square lattice consists of a d-wave flux phase for δ < δ_{c}, where δ is the doping away form half-filling, and a d-wave superconducting phase with a maximum T_{c} near δ_{c}. For δ < δ_{c} the flux phase lies above the superconducting phase and can be considered as a pseudogap phase. In order to confront the above phase diagram with experiment we have calculated the three irreducible electronic Raman spectra of A_{1g}, B_{1g}, and B_{2g} symmetry as a function of doping in the presence of superconductivity and the flux phase. We find that the B_{1g} spectrum is determined by transitions over the total gap with the main peak always increasing in energy with decreasing doping. For the A_{1g} spectrum we find that the indirect coupling to the superconducting order parameter is the leading one because it is unscreened if particle-hole symmetry holds. As a consequence the peak in the A_{1g} spectrum is caused by amplitude fluctuations of the superconducting order parameter, it is tied to T_{c} and decreases with decreasing δ below δ_{c}, similar as in the case of the B_{1g} spectrum. The good agreement of these results with experiment also supports the applicability of the above phase diagram to real high-T_{c} superconductors.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : R. Zeyher

Address for correspondence : R. Zeyher, Max-Planck-Institut für Festkörperforschung, Heisenbergstr. 1, 70506 Stuttgart, Germany

Email address : zeyher@audrey.mpi-stuttgart.mpg.de